OpenStudy (anonymous):

Simplify the difference quotient for the function given below, if x does not equal a. f(x)-f(a)/x-a for f(x)=sqrt(x-31)

OpenStudy (anonymous):

\[ \frac{\left(\sqrt{x-31}-\sqrt{a-31}\right) \left(\sqrt{a-31}-\sqrt{x-31}\right)}{(x-a) \left(\sqrt{a-31}-\sqrt{x-31}\right)} \]

OpenStudy (anonymous):

\[ \frac{\left(\sqrt{x-31}-\sqrt{a-31}\right) \left(\sqrt{a-31}-\sqrt{x-31}\right)}{(x-a) \left(\sqrt{a-31}-\sqrt{x-31}\right)}=\\\frac{x-a}{(x-a ) \left(\sqrt{a-31}-\sqrt{x-31}\right)}=\frac{1}{\sqrt{ a-31}-\sqrt{x-31}} \]

OpenStudy (anonymous):

It's \[\frac{ 1 }{ \sqrt{x-31}+\sqrt{a-31} }\] Thank you though!